OpenStudy (anonymous):

simplify. write your answer without parentheses. (x^2/-3y^3)^4

Nnesha (nnesha):

exponent rule:\[\huge\rm (a^m)^n = a^{m \times n}\] here is an example \[\huge\rm (2a^2)^3 =2^3 a^{3 \times 2}=8a^6\] both numbers/variables in the parentheses are raising to the outside power which is 3 in this examples